In a physics laboratory experiment, a coil with 250 turns enclosing an area of 12.1 cm2 is rotated in a time interval of 4.40×10−2 s from a position where its plane is perpendicular to the earth's magnetic field to one where its plane is parallel to the field. The earth's magnetic field at the lab location is 5.3×10−5 T .

Part A

What is the total magnetic flux through the coil before it is rotated?

Φ1=

Part B

What is the total magnetic flux through the coil after it is rotated?

Φ2=

Part C

What is the average emf induced in the coil?

E=
Answer #1

Part A -

Earth's magnetic field, B = 5.3 x 10^-5 T

Area of the coil, A = 12.1 cm^2 = 12.1 x 10^-4 m^2

No. of turns in the coil, N = 250

Total magnetic flux through the coil before it is rotated -

= N*B*A

= 250 *(5.3 x 10^-5) * (12.1 x 10^-4) = 1.603 x 10^-5 Wb

Part B -

Total magnetic flux through the coil after it is rotated -

= 0

Because, the coil has its plane parallel to the field.

Part C -

Average emf induced in the coil, E = change in flux/ time taken

= ( - ) / t

= (1.603 x 10^-5 - 0) / (4.40 x 10^-2) = 0.364 mV (Answer)